OpenStudy (anonymous):

how many solutions does x = y + 1 y = x + 1 have

OpenStudy (anonymous):

These are both equations for a line. Two lines will have either 0, 1, or infinitely many solutions. If the lines are parallel (same slope, different intercepts) there are 0 solutions. If the lines are the same (same slope, same intercept) there are infinite solutions. If they are different and not parallel (different slope) there is 1 solution. The two equations are y=x+1 y=x-1 So, based on this, how many solutions are there?